A family currently live in an apartment whose monthly rent is $950. They are thinking of buying a house which would cost $220,000. They plan to live in this house for 5 years and sell it at the end of the 5th year. They would put a downpayment of $20,000 and finance the balance through a mortgage at 3.5% interest rate. The mortgage is to be repaid in 5 annual installments (which include both principal and interest) at the end of each year for the next 5 years The house will have the following additional expenses: annual maintenance: $1500; Property taxes:$5500; Insurance: $1200. Assume they are in tax bracket of 20% and the price of home, rent and expenditure increases by 2.5% per year. Their opportunity cost or required rate of return is 5% per year. Note that property taxes are tax deductible and there no tax payable on capital gains. Use annual compounding for amortization schedule of mortgage.

Should they buy this house or continue to rent?

Buy since IRR is 5.47%

Buy since IRR is 5.51%

Rent since IRR is 4.16%
Buy since IRR is 6.28%

please show work

Thank you
Answer #1
Annuity on the mortgage
200000=Pmt.*(1-1.035^-5)/0.035
44296.27
1 2 3=Prev.5*3.5% 4=2-3 5=Prev.5-4 6=3*20%
Year Annuity Tow.Int. Tow.Mortg. Mortg. Bal. Int.tax shield
0 200000
1 44296.27 7000 37296.27 162703.7 1400
2 44296.27 5694.63 38601.64 124102.09 1138.93
3 44296.27 4343.57 39952.70 84149.39 868.71
4 44296.27 2945.23 41351.04 42798.35 589.05
5 44296.27 1497.94 42798.33 0.02 299.59
221481.35 21481.37 199999.98 4296.27
0 1 2 3 4 5
Cost -220000 248909.8
Opportunity cost of after-tax int.income lost on down pmt.(20000*5%*(1-20%) -800
After-tax Maintenance(1500*(-20%)) -1200 -1230 -1260.75 -1292.27 -1324.575
After-tax Property taxes(5500*1-20%)) -4400 -4510 -4622.75 -4738.32 -4856.777
After- tax insurance(1200*(1-20%)) -960 -984 -1008.6 -1033.82 -1059.66
After-tax savings in annual rent(950*12)*(1-20%) 9120 9348 9581.7 9821.243 10066.77
Int. tax shields(as per amortsn. Table) 1400 1138.93 868.71 589.05 299.59
Total annual cash flows -220800 3960 3762.93 3558.31 3345.89 252035.16
IRR 3.97%
Decision:
Rent since IRR 3.97% < reqd. return 5%